OpenStudy (anonymous):

proof? sinx+cosx=√2 sin(x+pi/4)

OpenStudy (anonymous):

\[\sin(x+\frac{\pi}{4})=\sin{x}\cos{\frac{\pi}{4}}+\sin{\frac{\pi}{4}}\cos{x}\]

OpenStudy (ash2326):

We have sin x+ cos x multiply the numerator and denominator by \(\sqrt 2\) \[ \frac{\sqrt 2}{\sqrt 2} (\sin x+ \cos x)\] We have now \[\frac{\sqrt 2}{1}(\sin x \frac{1}{\sqrt 2}+\cos x \frac{1}{\sqrt 2})\] now \[ \sin \pi/4= \cos \pi/4= \frac{1}{\sqrt 2}\] so we have \[\sqrt 2( \sin x \cos \pi/4+ \sin \pi/4 \cos x)\] we know \[ \sin (A+B)=\sin A \cos A+ \sin B \cos B\] so we get \[\sqrt 2( \sin (x +\pi/4))\]

OpenStudy (ash2326):

sorry \[\sin (A+B)=\sin A \cos B+ \sin B \cos A\]

OpenStudy (anonymous):

Hey ash2326 \[\sin(\pi/4)=\sqrt{2}/2\]

OpenStudy (anonymous):

\[\sin{(x+\frac{\pi}{4})}=\frac{1}{\sqrt{2}}(\sin{x}+\cos{x}) \]

OpenStudy (ash2326):

\[\sin \pi /4 =\sqrt 2 / 2 = 1/\sqrt 2\]

OpenStudy (anonymous):

and you got the demostrations in two ways. then it's done. =).

OpenStudy (anonymous):

ash began from the right to left, and I did from left to right.

OpenStudy (anonymous):

I proved it true? =\[\sqrt{2}\cos(x-\pi/4)\] =\[\sqrt{2}(cosx.\cos \pi/4 + sinx.\sin \pi/4)\] \[=cosx + sinx\] prove it?

OpenStudy (ash2326):

Yeah, this is also true

OpenStudy (anonymous):

Hooray for me

OpenStudy (ash2326):

Yeah, you did great:D

OpenStudy (anonymous):

متشکرم thanks

OpenStudy (anonymous):

hey I'm From Iran

OpenStudy (ash2326):

@hamidaarab Welcome:D

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

mr.ash2326 Do you know physics?

OpenStudy (anonymous):

find domain & Range of Function? f(x)=2sin^-1 (x+1)

OpenStudy (ash2326):

@hamidaarab don't have to thank so much. We all are friends here. I'll try but you should post your question in Physics group.

OpenStudy (anonymous):

How? Indeed ,find domain & Range of Function? f(x)=2sin^-1 (x+1)

OpenStudy (anonymous):

may you solve it?

OpenStudy (ash2326):

Yeah we have \[ f(x) = 2\sin^{-1} (x+1)\] \[ domain\ of\ \sin^{-1} x\ is\ [-1, 1]\] \[and\ range\ is (-90)\ degrees\ to\ (90)]\ degrees\] so domain of (x+1) is [-1,1] or domain of x is [-2,0] range of \( \sin^{-1} x\) is [-90, 90] or \([-\pi/2, \pi/2]\) so range of \( 2\sin^{-1} x\) is [-180,180] or \([-\pi, \pi]\)

OpenStudy (anonymous):

thank you very much ash2326